OpenStudy (girlgoyleh.):

What is the slope of the graph of 4x – 16y = 15

OpenStudy (anonymous):

change the equation in y=mx+c here m= slope

OpenStudy (girlgoyleh.):

is it 1/4?

OpenStudy (anonymous):

yes !!

OpenStudy (girlgoyleh.):

Thanks! By the way, do you know what an independent and dependent variable is?

OpenStudy (anonymous):

yes , I have an idea

OpenStudy (girlgoyleh.):

Could you explain it to me?

OpenStudy (anonymous):

1 moment plz

OpenStudy (girlgoyleh.):

Yep, thanks

OpenStudy (anonymous):

As an example suppose the function is the equation y = x^2 + 3 y depends on x through the function x^2 + 3. So for any specific value of x (the input) you get a value for y (the output). So for x = 1, say y = 1^2 +3 = 4 For x = 5, y = 5^2 +3 = 28 so x is independent variable here and y is dependent variable here

OpenStudy (girlgoyleh.):

Could you help me with another question?